# Data Guard: A Fine-grained Purpose-based Access Control System for Large Data Warehouses

### v1

@@ Abstract.
The last few years have witnessed a spate of data protection regulations in conjunction with an ever-growing appetite for data usage in large businesses, thus presenting significant challenges for businesses to maintain compliance. To address this conflict, we present Data Guard - a fine-grained, purpose-based access control system for large data warehouses. Data Guard enables authoring policies based on semantic descriptions of data and purpose of data access. Data Guard then translates these policies into SQL views that mask data from the underlying warehouse tables. At access time, Data Guard ensures compliance by transparently routing each table access to the appropriate data-masking view based on the purpose of the access, thus minimizing the effort of adopting Data Guard in existing applications. Our enforcement solution allows masking data at much finer granularities than what traditional solutions allow. In addition to row and column level data masking, Data Guard can mask data at the sub-cell level for columns with non-atomic data types such as structs, arrays, and maps. This fine-grained masking allows Data Guard to preserve data utility for consumers while ensuring compliance. We implemented a number of performance optimizations to minimize the overhead of data masking operations. We perform numerous experiments to identify the key factors that influence the data masking overhead and demonstrate the efficiency of our implementation.
PVLDB Reference Format:
Khai Tran, Sudarshan Vasudevan, Pratham Desai, Alex Gorelik, Mayank Ahuja, Athrey Yadatore Venkateshababu, Mohit Verma, Dichao Hu, Walaa Eldin Moustafa, Vasanth Rajamani, Ankit Gupta, Issac Buenrostro, Kalinda Raina. PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX Equal contributors.
Work done while the author was at LinkedIn Coporation.
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts are available at https://github.com/linkedin/dataguard-udfs.
1. Introduction
@@
The last two decades have witnessed a tremendous growth both in the collection and usage of data to power use cases such as business analytics, recommender systems and personalization. A significant amount of data that businesses collect from end users to power these applications is sensitive or personally identifiable. Understandably, there have been concerns from regulators, privacy advocates and various other citizen groups alike around the lack of transparency in the data collection and data handling practices adopted by businesses. Regulations such as GDPR (GDPR, 2016), CCPA (CCPA, 2018), HIPAA (HIPAA, 2003), DMA (DMA, 2024) and EU AI Act (EU AI Act, 2024) have been enacted into law in response to these privacy concerns and to drive transparency into the data collection and data usage practices of businesses. Specifically, these regulations require businesses to limit the usage of user data only for the purposes for which it is collected. As an example, the GDPR regulation requires businesses to seek consent from their end users (or formally, data subjects) in order to use sensitive attributes such as age and gender for advertising purposes. At the same time, the use of this data is often permitted in anonymized form to enable business use cases intended for demographic analysis.
@@
To address these privacy concerns and ensure compliance with regulations, several solutions have been proposed (Byun et al., 2005; Colombo and Ferrari, 2017; Xue et al., 2023) to enforce access control policies based on the purpose of data access. The systems are commonly referred to as Purpose-based Access Control (PBAC) systems. Depending on the granularity of enforcement, these systems can either authorize/deny access to an entire dataset (coarse-grained access control), or they can apply dynamic masking to hide contents of individual rows, columns or cells within the dataset (fine-grained access control). Enforcing policies dictated by these regulations across a large organization can be a challenging endeavor for the following reasons:
  1. (1)
@@
    A typical modern data stack involves multiple data processing systems such as Apache Spark (Spark, [n.d.]), Trino (Trino, [n.d.]) and Apache Flink (Flink, [n.d.]). Thus, any access control solution is required to apply data access policies consistently across a diversity of data processing systems used in an organization.
  2. (2)
@@
  3. (3)
    The data models underlying these datasets can be varied and complex. For instance, it is common to model data using non-atomic data types like arrays, maps, and structs. Furthermore, it is also common for AI/ML workloads to employ application-specific descriptions of data overlaid on basic data structures such as arrays and structs. A practical solution for access control must support granular data-masking capabilities within such complex data structures to maximize data utility while ensuring compliance.
Existing solutions have limited fine-grained access control capabilities and fall short in addressing one or more of the aforementioned challenges. Furthermore, we are not aware of any prior literature describing the myriad practical challenges encountered when deploying an access-control solution across a large data warehouse.
There are additional challenges for data-centric applications (such as LinkedIn among others) that employ a consent framework that controls how end-user data is used for different use cases. For instance, LinkedIn provides granular consent settings to its over one billion users (or data subjects) giving them control on how their data is used by LinkedIn for different use cases. Thus, data present in warehouse tables need to be dynamically masked based on data subject preferences in addition to the purpose of access. In Figure 1, we illustrate how access control based on data subject preferences differs from the more common row and column-level access control used in practice. In this example, the member_profiles table storing data subject information is queried for two different use cases, viz. “Ads” and “Jobs”. Data subject preferences controlling usage of different attributes of their data are stored in the member_settings table. Usage of data subject information for the Ads use case is controlled by two settings, allowEducationForAds and allowEmployerForAds, while that for Jobs use case is controlled by the settings allowEducationForJobs and allowEmployerForJobs. Based on the setting values for allowAgeEducationForAds and allowEmployerForAds, the education attribute associated with $memberId=123$ and employer attribute corresponding to $memberId=234$ in the member_profiles table need to be masked for the Ads use case. Similarly, the education attribute of member $234$ and the employer attribute of members $123$ need to be masked for Jobs use case based on the corresponding values for jobs-related settings.
id allowEdu ForAds allowEmpl ForAds allowEdu ForJobs allowEmpl ForJobs
@@
Figure 1. An example highlighting dynamic masking of data yielding different outputs based on purpose. The access control layer enforces member preferences (from the settings table) at data access time and selectively masks columns in the output.
To address these challenges, we designed and implemented Data Guard, a fine-grained purpose-based access control system for large data warehouses. The system translates policies specified in a domain-specific language (DSL) into purpose-specific views that mask data at access time. This translation is facilitated by maintaining a taxonomy of data classification labels (henceforth, referred to as policy labels), which are used to provide semantic description of the data elements (e.g. rows, columns, and cells in a table) present in the warehouse. The policies use purpose and policy labels to define the conditions under which access to a data element described using an annotation is allowed or denied. A policy compiler translates these rules into a purpose-specific view SQL and registered with the warehouse. At access time, a view routing layer transparently routes accesses to the appropriate data-masking view based on the purpose attached to the access.
@@
  • To the best of our knowledge, Data Guard is the first practical implementation of a fine-grained, purpose-based access control system that has been successfully deployed across a large enterprise warehouse. Data Guard has been rolled out at scale across LinkedIn’s warehouse ensuring all data accesses using Spark and Trino – the data processing engines used to query the warehouse data at LinkedIn – are compliant with our data usage policies.
  • Data Guard introduces a number of innovations that ensure both compliance and ease of use for warehouse users and application developers such as: (i) High-level abstractions that allow policies to be defined using purposes and semantic data labels. These policies are translated by Data Guard into data-masking views that serve as the data access API. (ii) A routing layer referred to as ViewShift which transparently routes accesses to warehouse tables from user queries to the appropriate data-masking views provisioned by Data Guard. (iii) View optimizations to minimize performance overhead in existing data processing jobs, such as the use of memory-efficient bitmaps in lieu of expensive table joins. (iv) Use of engine-agnostic views which allow data-masking views to be executable across different data processing engines (Spark and Trino at LinkedIn).
  • To the best of our knowledge, Data Guard is the first practical implementation of an access control system that allows data-masking at the sub-cell level. Data Guard achieves this via an expressive grammar that allows it to locate and mask deeply nested data elements inside composite attributes. These granular masking capabilities of Data Guard allows data consumers to maximize value from the underlying data while still ensuring compliance.
@@ 2. Related Work
Fine-grained access control in database management systems (DBMS) has been a well-studied topic and numerous solutions (Stonebraker and Wong, 1974; Rizvi et al., 2004; Xue et al., 2023; Agrawal et al., 2005) have been proposed to tackle this problem. In (Stonebraker and Wong, 1974), the authors propose a solution for the INGRES relational database that dynamically modifies user queries based on access restrictions defined on attributes in a relation. In (Xue et al., 2023), the authors propose a solution for Apache Spark applications that adds an access control enforcement stage in Spark’s Catalyst optimizer. This stage dynamically modifies non-compliant user queries into compliant ones. Unlike (Stonebraker and Wong, 1974; Xue et al., 2023),  (Rizvi et al., 2004) proposes using static authorization views which encode data access policies and admits a user query only if it can be mapped to an existing authorization view. A key limitation of these solutions is that they do not generalize to modern data stacks where the data models can be non-relational or can involve a diverse set of data processing engines.
The work proposed in (Agrawal et al., 2005) shares several advantages with Data Guard’s design including: (i) its ability to define policies that use semantic descriptions of data, and (ii) the ability to apply user preferences to dynamically filter tuples and attributes in a dataset. However, unlike Data Guard, the system described in (Agrawal et al., 2005) modifies user queries via dynamically created views. Indeed, we made an intentional design choice in Data Guard to create static versioned views to ensure query reproducibility and debuggability.
Several modern data warehouses such as Snowflake (Snowflake, 2024),
Databricks (Databricks, 2024) and Amazon Redshift (Amazon Redshift, 2024) provide capabilities to dynamically mask data based on semantic tags associated with attributes. However, they suffer from the same drawbacks as (Agrawal et al., 2005) in their inability to support query reproducibility and debuggability. Further, none of these solutions offer the scalability needed to support policies that require large joins with auxiliary tables such as the user consent tables used in our environment. In addition, these systems do not support data masking at the sub-cell level, a critical capability that Data Guard offers in order to maximize data utility.
Purpose-based access control for relational databases have been proposed in (Byun et al., 2005; Bertino, 2005) and further extended to non-relational data in (Byun and Li, 2008). These solutions describe a mechanism for associating an intended purpose with a data object (a relation, a tuple or an attribute) that is co-located with data. Further, they provide an algorithm to modify user query to filter data based on a comparison check between the intended purpose associated with the data object and the data accessor’s purpose. However, the proposed solutions only allow for filtering at a tuple-level granularity and do not support columnar or sub-cell-level filtering of data. Further, a major disadvantage of these solutions is that they require data to be rewritten when purposes change, which is impractical in large data warehouses where such changes are common.
There have been numerous works (Cranor et al., 2002; Ni et al., 2008; XACML, 2013) that provide policy language specifications for access control. The policy language used in Data Guard is similar in nature to the P3P policy language specification in (Cranor et al., 2002) in terms of its ability to use semantic tags to describe data items for which a policy is being defined. However, none of these solutions provide off-the-shelf language parsers and compilers required to translate the policies into machine representations for policy enforcement. Further, it was more cost-effective for us to define a constrained policy language and build a policy parser that fits our current needs.
Finally, there have been a number of works addressing privacy in databases e.g. (Agrawal et al., 2002; Chawla et al., 2005; Dwork, 2006).  (Agrawal et al., 2002) proposes a high-level system design for a privacy enforcing database and serves as a foundation for the work in (Agrawal et al., 2005).  (Dwork, 2006) introduces the notion of differential privacy and provides a mathematical definition for privacy loss in the query output.  (Chawla et al., 2005) provides a theoretical study of different metrics for measuring privacy in statistical databases and studies candidate techniques for privatization based on these metrics. While none of these works address the problem of access control studied in this paper, we believe Data Guard’s design is flexible enough to accommodate enforcement of policies that require privatization of query results.
3. Architecture Overview
Figure 2. Data Guard System Architecture.
@@ 3. Architecture Overview
Figure 2 shows the overall system architecture of Data Guard. Data producers provide semantic descriptions of data in the form of policy labels that are assigned to fields in a table. Fields refer to data elements in a table and represent the unit at which data masking is applied. Thus, a field can reference a column, a row, a cell or portion of data within a cell. The mapping of policy labels to fields is stored in a Data Catalog as shown in Figure 2. Policies in Data Guard are authored in a domain-specific language (DSL) and stored in a Policy Catalog which provides APIs to query policies from the catalog. Each policy is assigned to a purpose and policy label pair. A policy definition has an associated boolean condition that determines whether a data element with the corresponding policy label is masked or preserved during data access. Figure 2 shows an example policy that is assigned to $ads$ purpose and $employer$ label. The Policy Compiler consumes the policy label-field mapping and policy definitions as inputs and generates purpose-specific static views which are registered with the data warehouse. User queries submitted via Spark or Trino that access tables (e.g. $T_{1}$ and $T_{2}$ in Figure 2) are transparently redirected by a routing module called ViewShift to the appropriate data-masking views ($jobs.V_{1}$ and $ads.V_{2}$ respectively in Figure 2) based on the accessor’s purpose (i.e $jobs$ and $ads$ in Figure 2). We next describe the access control model and system implementation details of Data Guard architecture.
4. Access Control Model
@@
In this section, we introduce the main abstractions that enable purpose-limitation and fine-grained access control capabilities of Data Guard: (i) data access policies, (ii) data-masking views, and (iii) field paths.
4.1. Data Access Policies

### v2

@@ Abstract
The last few years have witnessed a spate of data protection regulations in conjunction with an ever-growing appetite for data usage in large businesses, which presents significant challenges for businesses to maintain compliance. To address this conflict, we present Data Guard - a fine-grained, purpose-based access control system for large data warehouses. Data Guard enables authoring policies based on semantic descriptions of data and purpose of data access. Data Guard then translates these policies into SQL views that mask data from the underlying warehouse tables. At access time, Data Guard ensures compliance by transparently routing each table access to the appropriate data-masking view based on the purpose of the access, thus minimizing the effort of adopting Data Guard in existing applications. Our enforcement solution allows masking data at much finer granularities than what traditional solutions allow. In addition to row and column level data masking, Data Guard can mask data at the sub-cell level for columns with non-atomic data types such as structs, arrays, and maps. This fine-grained masking allows Data Guard to preserve data utility for consumers while ensuring compliance. We implemented a number of performance optimizations to minimize the overhead of data masking operations. We perform numerous experiments to identify the key factors that influence the data masking overhead and demonstrate the efficiency of our implementation. Data Guard is deployed inside LinkedIn’s production data warehouses and ensures compliance of more than 20,000 table accesses each day across different data processing engines.
I Introduction
@@
The last two decades have witnessed a tremendous growth both in the collection and usage of data to power use cases such as business analytics, recommender systems and personalization. A significant amount of data that businesses collect from end users to power these applications, is sensitive or personally identifiable. Understandably, there have been concerns from regulators, privacy advocates and various other citizen groups alike concerning the lack of transparency in the data collection and data handling practices adopted by businesses. Regulations such as GDPR [1], CCPA [2], HIPAA [3], DMA [4] and EU AI Act [5] have been enacted into law in response to these privacy concerns and to drive transparency into the data collection and data usage practices of businesses. Specifically, these regulations require businesses to limit the usage of user data only for the purposes for which it is collected. As an example, the GDPR regulation requires businesses to seek consent from their end users (or formally, data subjects) in order to use sensitive attributes such as age and gender for advertising purposes. At the same time, the use of this data is often permitted in anonymized form to enable business use cases intended for demographic analysis.
@@
To address these privacy concerns and ensure compliance with regulations, several solutions have been proposed [6, 7, 8] to enforce access control policies based on the purpose of data access. The systems are commonly referred to as Purpose-based Access Control (PBAC) systems. Depending on the granularity of enforcement, these systems can either authorize/deny access to an entire dataset (coarse-grained access control), or they can apply dynamic masking to hide contents of individual rows, columns or cells within the dataset (fine-grained access control). Enforcing policies dictated by these regulations across a large organization can be a challenging endeavor for the following reasons:
  1. 1.
@@
    A typical modern data stack involves multiple data processing systems such as Apache Spark [9], Trino [10] and Apache Flink [11]. Thus, any access control solution is required to apply data access policies consistently across a diversity of data processing systems used in an organization.
  2. 2.
@@
  3. 3.
    The data models underlying these datasets can be varied and complex. For instance, it is common to model data using non-atomic data types like arrays, maps, and structs. Furthermore, it is also common for AI/ML workloads to employ application-specific descriptions of data overlaid on basic data structures such as arrays and structs. A practical solution for access control must support granular data-masking capabilities (at a sub-cell level) within such complex data structures to maximize data utility while ensuring compliance.
There are additional challenges for data-centric applications (such as LinkedIn among others) that employ a consent framework that controls how end-user data is used for different use cases. For instance, LinkedIn provides granular consent settings to its over one billion users (or data subjects) giving them control on how their data is used by LinkedIn for different use cases. Thus, data present in warehouse tables need to be dynamically masked based on data subject preferences in addition to the purpose of access. In Figure 1, we illustrate how access control based on data subject preferences differs from the more common row and column-level access control used in practice. In this example, the member_profiles table storing data subject information is queried for two different use cases, viz. “Ads” and “Jobs”. Data subject preferences controlling usage of different attributes of their data are stored in the member_settings table. Usage of data subject information for the Ads use case is controlled by two settings, allowEduForAds and allowEmpForAds, while that for Jobs use case is controlled by the settings allowEduForJobs and allowEmpForJobs. Based on the setting values for allowEduForAds and allowEmpForAds, the education attribute associated with $memberId=123$ and employer attribute corresponding to $memberId=234$ in the member_profiles table need to be masked for the Ads use case. Similarly, the education attribute of member $234$ and the employer attribute of members $123$ need to be masked for Jobs use case based on the corresponding values for jobs-related settings.
id allowEdu ForAds allowEmp ForAds allowEdu ForJobs allowEmp ForJobs
@@
Figure 1: An example highlighting dynamic masking of data to honor member preferences based on purpose yielding different outputs.
Existing solutions have limited fine-grained access control capabilities and fall short in addressing one or more of the aforementioned challenges. Furthermore, we are not aware of any prior literature describing the myriad practical challenges encountered when deploying an access-control solution across a large data warehouse supporting a diversity of data processing systems. We therefore implemented Data Guard, a fine-grained purpose-based access control system for large data warehouses. The system translates policies specified in a domain-specific language (DSL) into purpose-specific views that mask data at access time. This translation is facilitated by maintaining a taxonomy of data classification labels (henceforth, referred to as policy labels), which are used to provide semantic description of the data elements (e.g. rows, columns, and cells in a table) present in the warehouse. The policies use purpose and policy labels to define the conditions under which access to a data element described using an annotation is allowed or denied. A policy compiler translates these rules into a purpose-specific view SQL and registered with the warehouse. At access time, a view routing layer transparently routes accesses to the appropriate data-masking view based on the purpose attached to the access.
@@
  • Data Guard has been successfully deployed across LinkedIn’s exabyte-scale data warehouse and ensures that data accesses using Spark and Trino – the data processing engines used to query the warehouse data at LinkedIn – are compliant with our data usage policies, and honor preferences of >>1 billion LinkedIn members.
  • The implementation of Data Guard involves a number of novel ideas that ensure both compliance and ease of use for warehouse users and application developers such as: (i) High-level abstractions that allow policies to be defined using purposes and semantic data labels. These policies are translated by Data Guard into data-masking views that serve as the data access API. (ii) A routing layer referred to as ViewShift which transparently routes accesses to warehouse tables from user queries to the appropriate data-masking views provisioned by Data Guard. (iii) View optimizations to minimize performance overhead in existing data processing jobs, such as the use of memory-efficient bitmaps in lieu of expensive table joins. (iv) Use of engine-agnostic views which allow data-masking views to be executable across different data processing engines (Spark and Trino at LinkedIn).
  • To the best of our knowledge, Data Guard is the first practical implementation of an access control system that allows data masking at the sub-cell level. Data Guard achieves this via an expressive grammar that allows it to locate and mask deeply nested data elements inside composite attributes. These granular masking capabilities of Data Guard allows data consumers to maximize value from the underlying data while still ensuring compliance.
At LinkedIn, Data Guard is currently being used to enforce more than 110 unique data access policies for more than 50 purposes across two major production warehouses. Across the two major production warehouses, Data Guard actively maintains $\sim$13K views on $\sim$4K tables, having created more than 270K versioned views overall as labels and policies are updated. Each day, more than 20K table accesses across Spark and Trino are protected by Data Guard to ensure they are compliant with the defined data policies.
@@ II Related Work
Fine-grained access control in database management systems (DBMS) has been extensively studied, with numerous mechanisms proposed over the decades [12, 13, 8, 14, 15] that attempt to modify user’s data access queries into compliant ones. Early approaches such as [12], dynamically modified user queries to enforce attribute-level restrictions. More recently, Xue et. al. [8], propose a similar solution specifically for Spark applications by introducing an access control enforcement stage in Spark’s Catalyst optimizer.  [13] employs statically defined authorization views which encode data access policies and admits a user query only if it can be mapped to an existing authorization view. [15] enforces policies by dynamically rewriting queries to efficiently apply row-level masking. While effective in their original settings, these solutions fall short in modern data stacks, where access control often requires sub-cell-level masking and enforcement across heterogeneous processing engines.
Another line of work targets policy violation detection via static program analysis [16, 17, 18]. These approaches verify compliance but do not provide compliant query results at runtime. By contrast, Data Guard’s contribution is to automatically return compliant data to consumers, achieved through dynamic masking that incorporates both organizational policies and data-subject preferences.
The design in Agrawal et al. [14] shares several ideas with Data Guard, such as semantic descriptions of data and preference-aware filtering of tuples and attributes. However, their system relies on dynamically generated views to enforce restrictions. Further, the solution in [14] is described in the context of relational databases and does not therefore consider non-relational data models and sub-cell masking. In contrast, Data Guard employs static versioned views, a key technical choice that ensures query reproducibility and debuggability, both critical for production-scale analytics. Further, Data Guard is designed for modern data warehouses that allow non-relational data models and require sub-cell-level masking.
Commercial systems such as Snowflake [19], Databricks [20], and Amazon Redshift [21] support dynamic masking via attribute-level semantic tags. Yet, they inherit the drawbacks of [14]: lack of reproducibility, limited debuggability, and insufficient scalability for policies requiring large joins (e.g., user consent tables). Moreover, these systems generally cannot support sub-cell-level masking, which Data Guard introduces to maximize data utility while preserving compliance.
Purpose-based access control mechanisms proposed in [6, 22, 23] enrich relational models by associating intended usage purposes with data objects and filtering queries based on purpose compatibility. While conceptually elegant, these solutions are restricted to tuple-level enforcement, lack support for columnar or sub-cell masking, and require data rewriting when purposes evolve—which is impractical in large dynamic warehouses. Data Guard avoids such limitations by decoupling enforcement from data rewriting and supporting multiple levels of masking granularity.
Policy specification languages have also been explored in depth [24, 25, 26, 18]. Data Guard’s policy language draws inspiration from P3P [24] in its use of semantic tags to describe protected data. However, unlike these earlier efforts, we found no off-the-shelf compilers capable of translating such high-level specifications into enforcement-ready representations. Consequently, we designed a lightweight, constrained policy language and custom parser tailored for our current needs.
Finally, privacy-preserving databases [27, 28, 29] provide orthogonal but complementary mechanisms. For instance, Agrawal et al. [27] presented an early privacy-preserving DBMS architecture, while Dwork [29] introduced differential privacy, and Chawla et al. [28] formalized privacy metrics. While these works do not directly address access control, Data Guard’s flexible enforcement framework can incorporate such privacy-preserving techniques, extending its applicability to privacy-constrained analytics.
III Architecture Overview
Figure 2: Data Guard System Architecture.
@@ III Architecture Overview
Figure 2 shows the overall system architecture of Data Guard. Data producers provide semantic descriptions of data in the form of policy labels that are assigned to fields in a table. Fields refer to data elements in a table and represent the unit at which data masking is applied. Thus, a field can reference a column, a row, a cell or portion of data within a cell. The mapping of policy labels to fields is stored in a Data Catalog as shown in Figure 2. Policies in Data Guard are authored in a domain-specific language (DSL) and stored in a Policy Catalog which provides APIs to query policies from the catalog. Each policy is assigned to a purpose and policy label pair. A policy definition has an associated boolean condition that determines whether a data element with the corresponding policy label is masked or preserved during data access. Figure 2 shows an example policy that is assigned to adsads purpose and employeremployer label. The Policy Compiler consumes the policy label-field mapping and policy definitions as inputs and generates purpose-specific static views which are registered with the data warehouse. User queries submitted via Spark or Trino that access tables (e.g. $T_{1}$ and $T_{2}$ in Figure 2) are transparently redirected by a routing module called ViewShift to the appropriate data-masking views ($jobs.V_{1}$ and $ads.V_{2}$ respectively in Figure 2) based on the accessor’s purpose (i.e jobsjobs and adsads in Figure 2). We next describe the access control model and system implementation details of Data Guard architecture.
IV Access Control Model
@@
In this section, we introduce the foundational abstractions of Data Guard: (i) data access policies, (ii) data-masking views, and (iii) field paths. Together, these concepts enable the construction and deployment of a scalable access control system at LinkedIn that supports both fine-grained data protection and purpose limitation.
IV-A Data Access Policies